OpenStudy (anonymous):

Simplify the number using the imaginary unit i. sqrt[–16]

OpenStudy (anonymous):

\[i^2 = -1\] \[-1 = i^2 = \sqrt{-1}\sqrt{-1} = \sqrt{(-1)(-1)} = \sqrt{1} = 1\] \(i = \sqrt{-1}\) Are you able to continue from there?

OpenStudy (anonymous):

\(\sqrt{-16}\) \(\sqrt{16*(-1)} \) \(\sqrt{16} * \sqrt{-1} \) \((+/-) 4 * (+/-) i \) \((+/-) 4i \) @Apprentice am i right

OpenStudy (anonymous):

Yes, you are correct :-).